OpenStudy (anonymous):

When finding eigenvectors, what do you do when you reach a dead end like this? (example shown in comment)

OpenStudy (anonymous):

(-6 1) (0 -6) Lets say thats a matrix haha.....

OpenStudy (anonymous):

so solving for the eigenvalues, I got lamda(1,2)= -6 Then trying to make the eigenvector got me to 0n_1=n_2

OpenStudy (anonymous):

and i'm not sure how to proceed

OpenStudy (anonymous):

I got double root L =6. not -6. can you check and we work together?

OpenStudy (anonymous):

hmm....(-6-L)^2...wouldnt L=6 give you (-6-(6))^2=(-12)^2 I go tthat first too

OpenStudy (anonymous):

what is your characteristic equation? mine is L^2 -12L +36 = (L -6)^2 =0 ---> l =6

OpenStudy (anonymous):

oh, sorry, you are right, my bad.

OpenStudy (anonymous):

so, L = -6 and replace to matrix Ax -Lx

OpenStudy (anonymous):

wait no, I did what you said and now im torn.... If you find the DET you get (-6-L)(-6-L), which is (-6-L)^2.....(obviously haha) but if you multiply them together you get (L^2+12L+36), which could be turned into (L+6)^2, thus making you right......HHHOOOWW?

OpenStudy (anonymous):

OOOHH no nvm its -12L not +.....ok, so its L=-6....now what haha

OpenStudy (anonymous):

hahaaha... so weird . I am confused many times, when my discrete pro asks me using L-A while linear prof said that I must come up with A -L

OpenStudy (anonymous):

ok, anyway, it's yours, you check and choose which one is right, then we move.

OpenStudy (anonymous):

shoot i just got yours again....whatever, lets move on and Ill try both (I have unlimited attempts so its all good!)

OpenStudy (anonymous):

hi guy!! your brother must go somewhere to eat something, I cannot stay here to witness you switch many times like that. I'll be back in 15 minutes. do it carefully

OpenStudy (anonymous):

alright enjoy your meal, I got my final answer, and it's still L=-6. But I really just need to know how to go about turning that into and eigenvector and a generalized eigenvector. thanks for your help so far

OpenStudy (anonymous):

so, i guessed (1) (0) is my v_1 (0) (1) is my "generalized" eigenvector

OpenStudy (anonymous):

to count eigenvector, you use the matrix